Using Pythagorean Theorem, calculate the measure of b, if a = 4 cm. and c = 8 cm. In this case c is
the hypotenuse of the right angle triangle.

Answer by anantha(86)

sol:
in right angle triangle
(hypotenuse)^2=sum of the squares of the remaining two sides
this is Pythagorean Theorem
in this problem hypotenuse(c)=8cm and one side (a)=4cm
then we find the other side using above formula
c^2=a^2+b^2
8^2=4^2+b^2
64=16+b^2
subtract 16 on both sides
64-16=16-16+b^2
48=b^2
b= sqrt(48)
b=6.9 cm